# On sufficient conditions for the transitivity of homeomorphisms

### v3

@@ Abstract.
We derive a necessary and sufficient condition for a homeomorphism with the shadowing property to be topologically transitive: to have an invariant subset $A$, dense in the non-wandering set, where the barycenter property holds. To elucidate its dynamical nature, we compare this condition with other properties known to be sufficient for an Anosov diffeomorphism to be topologically transitive. We also describe the $C^{1}$ interior of the set of diffeomorphisms which comply with this condition, discuss examples with a variety of dynamics and present some applications of interest.
@@ 1. Introduction
Let $M$ be a $C^{\infty}$ closed (that is, connected, compact, boundaryless, finite dimensional) Riemannian manifold and ${\rm Diff}^{1}(M)$ denote the set of all $C^{1}$ diffeomorphisms of $M$ endowed with the $C^{1}$ topology. We say that $f\in{\rm Diff}^{1}(M)$ is Anosov if $M$ is a hyperbolic set for $f$, meaning that there are constants $c>0$ and $10<\lambda<$ 1 and a $Df-$invariant continuous splitting $T_{x}M=E^{s}_{f}(x)\oplus E^{u}_{f}(x)$ of the tangent space of $M$ at every $x\in M$ so that, for all non-negative integer $n$, one has
$\|Df^{n}_{x}(v)\|\,\leqslant\,c\,\lambda^{n}\,\|v\|\quad\forall v\in E^{s}_{f}(x)\quad\quad\text{and}\quad\quad\|Df^{-n}_{x}(v)\|\,\leqslant\,c\,\lambda^{n}\,\|v\|\quad\forall v\in E^{u}_{f}(x).$
@@
Let $(X,d)$ be a compact metric space, $T\colon X\to X$ be a homeomorphism and $A$ be a nonempty subset of $X$. We say that $T$ has the $su-$intersecting property in $A$ if $W^{s}_{T}(a)\cap W^{u}_{T}(b)\neq\emptyset$ for every $a,b\in A$. The homeomorphism $T$ is said to have the weak $su-$intersecting property in $A$ if $W^{s}_{T}\big{(}\overline{\mathcal{O}_{T}(a)}\big{)}\cap W^{u}_{T}\big{(}\overline{\mathcal{O}_{T}(b)}\big{)}\neq\emptyset$ for every $a,b\in A$.
Clearly, if an Anosov diffeomorphism $f\colon M\to M$ on a smooth closed Riemannian manifold $M$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$, then it is topologically transitive. Indeed, by Smale’s spectral decomposition theorem, the set $\Omega(f)$ is a finite union $\Omega_{1}\cup\cdots\cup\Omega_{K}$ of disjoint closed flimit-$f-$invariant pieces, the so called basic sets, on each of which $f$ is topologically transitive. Moreover, for each $1\leqslant j\leqslant K$, each $\Omega_{j}$ contains a dense subset of periodic orbits and $W^{s}_{f}(\Omega_{j})=\cup_{x\,\in\,\Omega_{j}}\,W^{s}_{f}(x)$, $W^{u}_{f}(\Omega_{j})=\cup_{x\,\in\,\Omega_{j}}\,W^{u}_{f}(x)$ and $M=\cup_{j=1}^{K}\,W^{s}_{f}(\Omega_{j})=\cup_{j=1}^{K}\,W^{u}_{f}(\Omega_{j})$ (cf. [11, Proposition 3.10]). Take a repeller $\Omega_{r}$ and an attractor $\Omega_{a}$, whose existence is due to the adapted filtration for the spectral decomposition of $\Omega(f)$ (cf. [55, § 2]). It is known that $\Omega_{r}$ contains the stable manifolds of all its points and, similarly, $\Omega_{a}$ contains the unstable manifolds of its elements (cf. [11, Lemma 4.9]). So, if $x\in\Omega_{r}\cap\operatorname{Per}(f)$ and $y\in\Omega_{a}\cap\operatorname{Per}(f)$, then $W^{s}_{f}(x)\subset\Omega_{r}$, $W^{u}_{f}(x)\subset\Omega_{a}$ and, by the weak $su-$intersecting property in $\operatorname{Per}(f)$, one also has $W^{s}_{f}(x)\cap W^{u}_{f}(y)\neq\emptyset$. Hence $\Omega_{r}\cap\Omega_{a}\neq\emptyset$, which is a contradiction unless $f$ has only one basic set equal to $M$. Thus $f$ is topologically transitive.
@@
Let $(X,d)$ be a compact metric space, $T\colon X\to X$ be a homeomorphism and $A\subseteq X$ be a nonempty set. We say that $T$ has the barycenter property in $A$ if for every $p,\,q\in A$ and any $\varepsilon>0$, there is a positive integer $N=N(\varepsilon,p,q)$ such that, for any $n_{1},n_{2}\in\mathbb{N}$, one can find an integer $m\in[0,N]$ and $x_{0}\in X$ satisfying $d\big{(}T^{i}(x_{0}),T^{i}(p)\big{)}<\varepsilon$ for every $-n_{1}\leqslant i\leqslant 0$ and $d\big{(}T^{i+m}(x_{0}),T^{i}(q)\big{)}<\varepsilon$ for every $0\leqslant i\leqslant n_{2}$.
It turns out that, if $T$ has the weak $su-$intersecting property in $\operatorname{Per}(T)$, then $T$ has the barycenter property in $\operatorname{Per}(T)$ (cf. [58, Lemma 2.2]). Moreover, if $X$ has no isolated points and $T$ is expansive, then $T$ has the weak $su-$intersecting property in $\operatorname{Per}(T)$ if and only if $T$ has the barycenter property in $\operatorname{Per}(T)$. Indeed, the proof of the aforementioned Lemma 2.2 in [58] only requires that for every periodic point $p$ of $T$ there exists $\varepsilon>0$ such that $W^{s}_{T,\,\varepsilon}(p)\subseteq W^{s}_{T}(p)$ and $W^{u}_{T,\,\varepsilon}(p)\subseteq W^{u}_{T}(p)$; and, surely, this condition is valid when $T$ is expansive (cf. [39, Lemma 1]). This ultimately implies that, within the setting of $C^{1}$ diffeomorphisms of a closed Riemannian manifold whose periodic points are hyperbolic, the barycenter property in the set of periodic points is equivalent to the weak $su-$intersecting property in that set. For instance, within the family of $C^{1}$ $\Omega-$stable diffeomorphisms we may equally use one or the other property.
@@
We observe that the assumption $\Omega(T)=CR(T)$ holds for homeomorphisms with the shadowing property [6, Theorem 3.1.2]; generically in ${\rm Diff}^{1}(M)$ [10, Corollaire 1.2]; and, by the Closing Lemma [49], for $\Omega$-stable diffeomorphisms if $M$ has dimension $\geqslant 2$. Furthermore, the condition $\Omega(T)=CR(T)$ is not equivalent to shadowing, as exemplified by an irrational rotation on the unit circle, which satisfies $\Omega(T)=CR(T)$ but does not exhibit shadowing.
@@
In [34], M. Lee proved that the $C^{1}$ interior of the diffeomorphisms satisfying the barycenter property in $\operatorname{Per}(f)$ is contained in the set of $\Omega-$stable diffeomorphisms; and that, generically, if a $C^{1}$ diffeomorphism has the barycenter property in $\operatorname{Per}(f)$ then it is $\Omega-$stable. We will improve this information by showing that the $C^{1}$ interior of diffeomorphisms satisfying the weak $su-$intersecting property in $\operatorname{Per}_{h}(f)$ is equal to the set of topologically transitive Anosov diffeomorphisms (see Proposition 7.10); and by proving that, generically, if a $C^{1}$ diffeomorphism has the weak $su-$intersecting property in $\operatorname{Per}_{h}(f)$, then it is a topologically transitive Anosov diffeomorphism (see Subsection 7.3.2).
@@
    $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$.
@@ 2.7.6. Specification
is an $N-$spaced specification if $a_{i}-b_{i-1}\geqslant N$ for every $2\leqslant i\leqslant n$. Given $\varepsilon>0$, we say that a specification $\xi$ is $\varepsilon-$traced by $y\in X$ if
[…]
The map $T$ has the specification property if for any $\varepsilon>0$ there is $N=N(\varepsilon)\in\mathbb{N}$ such that any $N-$spaced specification $\xi$ is $\varepsilon-$traced by some $y\in X$.
@@ 2.8. Gluing orbit property
[…]
@@ 2.9. Measure center
In what follows, $\mathcal{P}(X)$ stands for the space of Borel probability measures in $X$ with the ${}^{*}-$topology, $\mathcal{P}_{T}(X)$ is its subspace of $T-$invariant elements and $\mathcal{P}_{T}^{erg}(X)$ is the subset of ergodic measures of $\mathcal{P}_{T}(X)$. The measure center of $T$ is the set
@@ 3. Examples
Suppose that $X$ possesses a Lebesgue-like measure, say $\mu_{0}$. Denote by $\operatorname{Hom}(X,\mu_{0})$ the set of homeomorphisms of $X$ which preserve $\mu_{0}$. This space is metrizable by the uniform distance for homeomorphisms, given by $D(f,g)=D_{1}(f,g)+D_{1}(f^{-1},g^{-1})$, where $D_{1}(f,g)=\max_{x\,\in\,X}d(f(x),g(x))$. Moreover, $(\operatorname{Hom}(X,\mu_{0}),D)$ is a Baire space.
@@
and let $\Sigma_{A}\subset Y^{\mathbb{Z}}$ be the space of $A-$admissible sequences, that is,
[…]
@@
We observe that, on the contrary, if $T\colon X\to X$ is a homeomorphism of a compact metric space $(X,d)$ without isolated points, $A$ is a nonempty subset of $X$ and $T$ has the specification property, then $T$ has the barycenter property in $A$. Indeed, take $p,\,q\in A$ and fix $\varepsilon>0$. Since $T$ has the specification property there exists $N=N(\varepsilon)\in\mathbb{N}$ such that any $N-$spaced specification $\xi$ is $\varepsilon-$traced by some $y\in X$. Given $n_{1},n_{2}\in\mathbb{N}$, define
@@
is an $N-$spaced specification. Then $\xi$ is $\varepsilon-$traced by some $y\in X$, that is,
[…]
@@
Claim 2: The $su-$intersecting property is invariant by iteration.
@@
Claim 3: The $su-$intersecting property induces some rigidity on the dynamics.
@@
$W^{s}_{T}(1)$ $=$ $\big{\{}1\big{\}}\cup\big{\{}1/n\,\colon\,n\geqslant 2\big{\}}\cup\big{\{}1-1/n\,\colon\,n\geqslant 3\big{\}}\quad\quad\text{and}\quad\quad W^{u}_{T}(1)\,=\,\big{\{}1\big{\}}$
$W^{u}_{T}(0)$ $=$ $\big{\{}0\big{\}}\cup\big{\{}1/n\,\colon\,n\geqslant 2\big{\}}\cup\big{\{}1-1/n\,\colon\,n\geqslant 3\big{\}}\quad\quad\text{and}\quad\quad W^{s}_{T}(0)\,=\,\{0\}.$
So, $T$ does not have the $su-$intersecting property in $\operatorname{Per}(T)$. Moreover, if $p=1$ and $q=0$ then, given $20<\varepsilon<1/$ 2, the subset of the orbit of $1/2$ by $T$ outside the union $[0,\varepsilon[\,\cup\,\,]1-\varepsilon,1]$ is finite. Let
@@
Therefore, for any $n_{1},n_{2}\in\mathbb{N}$, one can find an integer $m=N\in[0,N]$ and $x_{0}\in X$ satisfying $d\big{(}T^{i}(x_{0}),T^{i}(p)\big{)}<\varepsilon$ for every $-n_{1}\leqslant i\leqslant 0$ and $d\big{(}T^{i+m}(x_{0}),T^{i}(q)\big{)}<\varepsilon$ for every $0\leqslant i\leqslant n_{2}$. Thus, the pair $p,q$ complies with the condition of the barycenter property in $\operatorname{Per}(T)$. However, if we exchange the fixed points this condition is no longer valid. Thus, $T$ does not have the barycenter property in $\operatorname{Per}(T)$. We note that $\Omega(T)=\overline{\operatorname{Per}(T)}$, $T$ has the shadowing property (see [46, Proposition 3]) and is topologically transitive. However, $X$ has isolated points and the conclusion of Theorem A(c) does not hold.
@@
The next example calls our attention to the relevance of the uniform hyperbolicity assumption in Theorem C. Let $f\colon M\to M$ be a $C^{1}$ Anosov diffeomorphism of the $2-$dimensional torus and $g\colon M\to M$ be a $C^{1}$ almost Anosov diffeomorphism obtained by H. Hu in [28] through a first perturbation on $f$. The resulting diffeomorphism $g$ is uniformly hyperbolic except at one fixed point $p$, where $Df_{p}$ is the identity map. The stable and unstable foliations of $f$ persist under the perturbation, as well as their intersections. So $g$ has the $su-$intersecting property in $\operatorname{Per}(g)$, inherited from $f$. By construction, $g$ is topologically transitive, but is not an Anosov diffeomorphism. A similar construction may be found in [27].
@@
Recall that, in [12], M. Brin proved that if $f\colon M\to M$ is a $C^{1}$ strong partially hyperbolic diffeomorphism of a closed Riemannian manifold such that $\Omega(f)=M$ and $f$ is accessible in $M$, then $f$ is topologically transitive. In order to show that Brin’s result cannot be generalized to the case where the chain recurrent set of $f$ is the whole manifold (that is, when $f$ is chain-transitive), S. Gan and Y. Shi constructed in [22] a partially hyperbolic diffeomorphism on the $3-$dimensional torus $\mathbb{T}^{3}$ which is accessible in $M$ and chain-transitive, but fails to be topologically transitive. Notice, however, that, as a chain-transitive diffeomorphism which has the shadowing property is topologically transitive [31, Lemma 2.1], the example of S. Gan and Y. Shi has no such property.

### v4

@@ Abstract.
We derive sufficient conditions for a homeomorphism of a compact metric space without isolated points to be topologically transitive. In particular, we prove that a homeomorphism with the shadowing property has a dense orbit if either it has the su-intersecting property in the whole space; or it has the barycenter property in the non-wandering set; or else it is accessible in the recurrent set. To elucidate their dynamical nature, we compare these three conditions, discuss examples with a variety of dynamics and present some applications of interest.
@@ 1. Introduction
Let $M$ be a $C^{\infty}$ closed (that is, connected, compact, boundaryless, finite dimensional) Riemannian manifold and ${\rm Diff}^{1}(M)$ denote the set of all $C^{1}$ diffeomorphisms of $M$ endowed with the $C^{1}$ topology. We say that $f\in{\rm Diff}^{1}(M)$ is Anosov if $M$ is a hyperbolic set for $f$, meaning that there are constants $c>0$ and $0<\lambda<1$ and a DfDf-invariant continuous splitting $T_{x}M=E^{s}_{f}(x)\oplus E^{u}_{f}(x)$ of the tangent space of $M$ at every $x\in M$ so that, for all non-negative integer $n$, one has
$\|Df^{n}_{x}(v)\|\,\leqslant\,c\,\lambda^{n}\,\|v\|\quad\forall v\in E^{s}_{f}(x)\quad\text{and}$
$\|Df^{-n}_{x}(v)\|\,\leqslant\,c\,\lambda^{n}\,\|v\|\quad\forall v\in E^{u}_{f}(x).$
@@
Let $(X,d)$ be a compact metric space, $T\colon X\to X$ be a homeomorphism and $A$ be a nonempty subset of $X$. We say that $T$ has the $su$-intersecting property in $A$ if $W^{s}_{T}(a)\cap W^{u}_{T}(b)\neq\emptyset$ for every $a,b\in A$. The homeomorphism $T$ is said to have the weak susu-intersecting property in $A$ if $W^{s}_{T}\big{(}\overline{\mathcal{O}_{T}(a)}\big{)}\cap W^{u}_{T}\big{(}\overline{\mathcal{O}_{T}(b)}\big{)}\neq\emptyset$ for every $a,b\in A$.
Clearly, if an Anosov diffeomorphism $f\colon M\to M$ on a smooth closed Riemannian manifold $M$ has the weak susu-intersecting property in $\operatorname{Per}(f)$, then it is topologically transitive. Indeed, by Smale’s spectral decomposition theorem, the set $\Omega(f)$ is a finite union $\Omega_{1}\cup\cdots\cup\Omega_{K}$ of disjoint closed $f$-invariant pieces, the so called basic sets, on each of which $f$ is topologically transitive. Moreover, for each $1\leqslant j\leqslant K$, each $\Omega_{j}$ contains a dense subset of periodic orbits and $W^{s}_{f}(\Omega_{j})=\cup_{x\,\in\,\Omega_{j}}\,W^{s}_{f}(x)$, $W^{u}_{f}(\Omega_{j})=\cup_{x\,\in\,\Omega_{j}}\,W^{u}_{f}(x)$ and $M=\cup_{j=1}^{K}\,W^{s}_{f}(\Omega_{j})=\cup_{j=1}^{K}\,W^{u}_{f}(\Omega_{j})$ (cf. [11, Proposition 3.10]). Take a repeller $\Omega_{r}$ and an attractor $\Omega_{a}$, whose existence is due to the adapted filtration for the spectral decomposition of $\Omega(f)$ (cf. [55, § 2]). It is known that $\Omega_{r}$ contains the stable manifolds of all its points and, similarly, $\Omega_{a}$ contains the unstable manifolds of its elements (cf. [11, Lemma 4.9]). So, if $x\in\Omega_{r}\cap\operatorname{Per}(f)$ and $y\in\Omega_{a}\cap\operatorname{Per}(f)$, then $W^{s}_{f}(x)\subset\Omega_{r}$, $W^{u}_{f}(x)\subset\Omega_{a}$ and, by the weak susu-intersecting property in $\operatorname{Per}(f)$, one also has $W^{s}_{f}(x)\cap W^{u}_{f}(y)\neq\emptyset$. Hence $\Omega_{r}\cap\Omega_{a}\neq\emptyset$, which is a contradiction unless $f$ has only one basic set equal to $M$. Thus $f$ is topologically transitive.
@@
Let $(X,d)$ be a compact metric space, $T\colon X\to X$ be a homeomorphism and $A\subseteq X$ be a nonempty set. We say that $T$ has the barycenter property in $A$ if for every $p,\,q\in A$ and any $\varepsilon>0$, there is a positive integer $N=N(\varepsilon,p,q)$ such that, for any $n_{1},n_{2}\in\mathbb{N}$, one can find an integer $m\in[0,N]$ and $x_{0}\in X$ satisfying $d\big{(}T^{i}(x_{0}),T^{i}(p)\big{)}<\varepsilon$ for every $n_{1}\leqslant i\leqslant 0$ and $d\big{(}T^{i+m}(x_{0}),T^{i}(q)\big{)}<\varepsilon$ for every $0\leqslant i\leqslant n_{2}$.
It turns out that, if $T$ has the weak susu-intersecting property in $\operatorname{Per}(T)$, then $T$ has the barycenter property in $\operatorname{Per}(T)$ (cf. [58, Lemma 2.2]). Moreover, if $X$ has no isolated points and $T$ is expansive, then $T$ has the weak susu-intersecting property in $\operatorname{Per}(T)$ if and only if $T$ has the barycenter property in $\operatorname{Per}(T)$. Indeed, the proof of the aforementioned Lemma 2.2 in [58] only requires that for every periodic point $p$ of $T$ there exists $\varepsilon>0$ such that $W^{s}_{T,\,\varepsilon}(p)\subseteq W^{s}_{T}(p)$ and $W^{u}_{T,\,\varepsilon}(p)\subseteq W^{u}_{T}(p)$; and, surely, this condition is valid when $T$ is expansive (cf. [39, Lemma 1]). This ultimately implies that, within the setting of $C^{1}$ diffeomorphisms of a closed Riemannian manifold whose periodic points are hyperbolic, the barycenter property in the set of periodic points is equivalent to the weak susu-intersecting property in that set. For instance, within the family of $C^{1}$ $\Omega$-stable diffeomorphisms we may equally use one or the other property.
@@
We observe that the assumption $\Omega(T)=CR(T)$ holds for homeomorphisms with shadowing property [6, Theorem 3.1.2]; generically in ${\rm Diff}^{1}(M)$ [10, Corollaire 1.2] and, by the Closing Lemma [49], for $\Omega$-stable diffeomorphisms if $M$ has dimension $\geqslant 2$. Furthermore, the condition $\Omega(T)=CR(T)$ is not equivalent to shadowing, as exemplified by an irrational rotation on the unit circle, which satisfies $\Omega(T)=CR(T)$ but does not exhibit shadowing.
@@
In [34], M. Lee proved that the $C^{1}$ interior of the diffeomorphisms satisfying the barycenter property in $\operatorname{Per}(f)$ is contained in the set of $\Omega$-stable diffeomorphisms; and that, generically, if a $C^{1}$ diffeomorphism has the barycenter property in $\operatorname{Per}(f)$ then it is $\Omega$-stable. We will improve this information by showing that the $C^{1}$ interior of diffeomorphisms satisfying the weak susu-intersecting property in $\operatorname{Per}_{h}(f)$ is equal to the set of topologically transitive Anosov diffeomorphisms (see Proposition 7.10); and by proving that, generically, if a $C^{1}$ diffeomorphism has the weak susu-intersecting property in $\operatorname{Per}_{h}(f)$, then it is a topologically transitive Anosov diffeomorphism (see Subsection 7.3.2).
@@
    $f$ has the weak susu-intersecting property in $\operatorname{Per}(f)$.
@@ 2.7.6. Specification
is an NN-spaced specification if $a_{i}-b_{i-1}\geqslant N$ for every $2\leqslant i\leqslant n$. Given $\varepsilon>0$, we say that a specification $\xi$ is $\varepsilon$-traced by $y\in X$ if
$d\big{(}T^{k}(y),T^{k}(x_{i})\big{)}\,<\,\varepsilon\quad\quad\forall\,a_{i}\,\leqslant\,k\,\leqslant\,b_{i}\quad\forall\,1\,\leqslant\,i\,\leqslant\,n.$
The map $T$ has the specification property if for any $\varepsilon>0$ there is $N=N(\varepsilon)\in\mathbb{N}$ such that any NN-spaced specification $\xi$ is $\varepsilon$-traced by some $y\in X$.
@@ 2.8. Gluing orbit property
$d\big{(}T^{j}(x),\,T^{j}(x_{1})\big{)}$ $\leqslant$ $\varepsilon\quad\quad\forall\,0\leqslant j\leqslant n_{1}$
$d\big{(}T^{j\,+\,n_{1}\,+\,p_{1}\,+\,\dots\,+\,n_{i-1}\,+\,p_{i-1}}(x),\;T^{j}(x_{i})\big{)}$ $\leqslant$ $\varepsilon\quad\quad\forall\,2\leqslant i\leqslant k\quad\forall\,0\leqslant j\leqslant n_{i}.$
@@ 2.9. Measure center
[…]
@@ 3. Examples
Suppose that $X$ possesses a Lebesgue-like measure, say $\mu_{0}$. Denote by $\operatorname{Hom}(X,\mu_{0})$ the set of homeomorphisms of $X$ which preserve $\mu_{0}$. This space is metrizable by the uniform distance for homeomorphisms, given by
$D(f,g)=D_{1}(f,g)+D_{1}(f^{-1},g^{-1}),$
where $D_{1}(f,g)=\max_{x\,\in\,X}d(f(x),g(x))$. Moreover, $(\operatorname{Hom}(X,\mu_{0}),D)$ is a Baire space.
@@
and let $\Sigma_{A}\subset Y^{\mathbb{Z}}$ be the space of AA-admissible sequences, that is,
$\Sigma_{A}\,=\,\big{\{}(x_{k})_{k\,\in\,\mathbb{Z}}\colon\,a_{x_{k}\,x_{k+1}}\,=\,1\quad\forall k\in\mathbb{Z}\big{\}}.$
@@
We observe that, on the contrary, if $T\colon X\to X$ is a homeomorphism of a compact metric space $(X,d)$ without isolated points, $A$ is a nonempty subset of $X$ and $T$ has the specification property, then $T$ has the barycenter property in $A$. Indeed, take $p,\,q\in A$ and fix $\varepsilon>0$. Since $T$ has the specification property there exists $N=N(\varepsilon)\in\mathbb{N}$ such that any NN-spaced specification $\xi$ is $\varepsilon$-traced by some $y\in X$. Given $n_{1},n_{2}\in\mathbb{N}$, define
@@
is an NN-spaced specification. Then $\xi$ is $\varepsilon$-traced by some $y\in X$, that is,
$d\big{(}T^{k}(y),\,T^{k}(p)\big{)}$ $<$ $\varepsilon\quad\quad\,-n_{1}\,=\,a_{1}\,\leqslant\,k\,\leqslant\,b_{1}\,=\,0$
@@
Claim 2: The susu-intersecting property is invariant by iteration.
@@
Claim 3: The susu-intersecting property induces some rigidity on the dynamics.
@@
$W^{s}_{T}(1)$ $=$ $\big{\{}1\big{\}}\cup\big{\{}1/n\,\colon\,n\geqslant 2\big{\}}\cup\big{\{}1-1/n\,\colon\,n\geqslant 3\big{\}},\quad W^{u}_{T}(1)\,=\,\big{\{}1\big{\}}$
$W^{u}_{T}(0)$ $=$ $\big{\{}0\big{\}}\cup\big{\{}1/n\,\colon\,n\geqslant 2\big{\}}\cup\big{\{}1-1/n\,\colon\,n\geqslant 3\big{\}},\quad W^{s}_{T}(0)\,=\,\big{\{}0\big{\}}.$
So, $T$ does not have the $su$-intersecting property in $\operatorname{Per}(T)$. Moreover, if $p=1$ and $q=0$ then, given $0<\varepsilon<1/2$, the subset of the orbit of $1/2$ by $T$ outside the union $0,\varepsilon[\,\cup\,\,]1-\varepsilon,1]$ is finite. Let
@@
Therefore, for any $n_{1},n_{2}\in\mathbb{N}$, one can find an integer $m=N\in[0,N]$ and $x_{0}\in X$ satisfying $d\big{(}T^{i}(x_{0}),T^{i}(p)\big{)}<\varepsilon$ for every $n_{1}\leqslant i\leqslant 0$ and $d\big{(}T^{i+m}(x_{0}),T^{i}(q)\big{)}<\varepsilon$ for every $0\leqslant i\leqslant n_{2}$. Thus, the pair $p,q$ complies with the condition of the barycenter property in $\operatorname{Per}(T)$. However, if we exchange the fixed points this condition is no longer valid. Thus, $T$ does not have the barycenter property in $\operatorname{Per}(T)$. We note that $\Omega(T)=\overline{\operatorname{Per}(T)}$, $T$ has the shadowing property (see [46, Proposition 3]) and is topologically transitive. However, $X$ has isolated points and the conclusion of Theorem A(c) does not hold.
@@
The next example calls our attention to the relevance of the uniform hyperbolicity assumption in Theorem C. Let $f\colon M\to M$ be a $C^{1}$ Anosov diffeomorphism of the 22-dimensional torus and $g\colon M\to M$ be a $C^{1}$ almost Anosov diffeomorphism obtained by H. Hu in [28] through a first perturbation on $f$. The resulting diffeomorphism $g$ is uniformly hyperbolic except at one fixed point $p$, where $Df_{p}$ is the identity map. The stable and unstable foliations of $f$ persist under the perturbation, as well as their intersections. So $g$ has the $su$-intersecting property in $\operatorname{Per}(g)$, inherited from $f$. By construction, $g$ is topologically transitive, but is not an Anosov diffeomorphism. A similar construction may be found in [27].
@@
Recall that, in [12], M. Brin proved that if $f\colon M\to M$ is a $C^{1}$ strong partially hyperbolic diffeomorphism of a closed Riemannian manifold such that $\Omega(f)=M$ and $f$ is accessible in $M$, then $f$ is topologically transitive. In order to show that Brin’s result cannot be generalized to the case where the chain recurrent set of $f$ is the whole manifold (that is, when $f$ is chain-transitive), S. Gan and Y. Shi constructed in [22] a partially hyperbolic diffeomorphism on the 33-dimensional torus $\mathbb{T}^{3}$ which is accessible in $M$ and chain-transitive, but fails to be topologically transitive. Notice, however, that, as a chain-transitive diffeomorphism which has the shadowing property is topologically transitive [31, Lemma 2.1], the example of S. Gan and Y. Shi has no such property.